# Joint moments of characteristic polynomials from the orthogonal and unitary symplectic groups

### v1

@@ 4.2.3 A recursive relation for $\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{n,j}}$
where $P_{m,1}^{(a,b)}(t)$ and $P_{m,2}^{(a,b)}(t)$ are polynomials of $t$. Moreover, the coefficients of these polynomials are themselves polynomials in $N,a$. Note that, $\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\mathrm{e}^{-t\mathfrak{p}_{1,N}}\right]$ is $\sum_{k=1}^{m}kn_{k}$-times differentiable and it can be represented as $\exp\left(\int_{0}^{t}\frac{\sigma_{N}(t)}{t}\mathrm{d}t\right)$, where $\sigma_{N}(t)$ satisfies $\sigma$-Painlevé-V equation (8). Hence, examining the asymptotic expansion near $t\sim 0$ (in powers of $t$) of both $\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\mathrm{e}^{-t\mathfrak{p}_{1,N}}\right]$ and $\sigma_{N}(t)$, where the latter can be obtained by iteratively using the $\sigma$-Painlevé-V equation (a similar argument to that of [14, Section 4.2]), we obtain that the Taylor coefficients of $\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\mathrm{e}^{-t\mathfrak{p}_{1,N}}\right]$ are of the form $\frac{\mathcal{Q}_{2}(N;a)}{\mathcal{Q}_{1}(a)}$, where $\mathcal{Q}_{1}(a)$ is a polynomial in $a$ and $\mathcal{Q}_{2}(N;a)$ is a polynomial in $a$, whose coefficients themselves are polynomials in $N$. After plugging this into the right-hand side of equation (4.4), while noting that the coefficients of the singular terms $\frac{1}{t^{l}}$ for $l=1,\dots,\sum_{k=1}^{m}kn_{k}$ vanish (which follows by again considering the asymptotic expansion at $t\sim 0$ and noting that it has to be bounded due to Theorem 1.9), we see that right-hand side of equation (4.4) must be a rational function in $a$ of the form described above; and therefore, so does the left-hand side of equation (4.4). Hence, by Newton’s formula, we readily conclude that $\mathbb{E}_{N}^{(a,-\frac{1}{2}{})}\left[\prod_{k=1}^{m}\mathfrak{e}_{N,k}^{h_{k}}\right]$ has the rational function form $\frac{Q_{2}(N;a)}{Q_{1}(a)}$, for certain polynomials $Q_{1}(a)$ and $Q_{2}(N;a)$ in $a$, whose coefficients of the powers of $a$ in $Q_{2}(N;a)$ are polynomials in $N$ whenever $\Re(a)>\sum_{k=1}^{m}kh_{k}-1$. On the other hand, the map $a\mapsto\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\prod_{k=1}^{m}\mathfrak{e}_{N,k}^{h_{k}}\right]$, which is initially defined on a part of the real line, extends trivially to an analytic function on the set $\{a\in\mathbb{C}\mid\Re(a)>\sum_{k=1}^{m}h_{k}-1\}$ Thus, by the identity theorem, it must equal $\frac{Q_{2}(N;a)}{Q_{1}(a)}$ whenever $\Re(a)>\sum_{k=1}^{m}h_{k}-1$. Since $\sum_{k=0}^{m}h_{k}-\frac{1}{2}>\sum_{k=1}^{m}h_{k}-1$ when $h_{0}\geq 0$, $\frac{Q_{2}(N;a)}{Q_{1}(a)}$ is a rational function of $h_{0}$ on all of $h_{0}\geq 0$. Note that, when $h_{0}>\frac{1}{2}$, $N^{-2\sum_{k=1}^{m}{kh_{k}}}\frac{Q_{2}(N;a)}{Q_{1}(a)}$ converges as $N\rightarrow\infty$. Since this is a rational function, this convergence holds for any $h_{0}\geq 0$ and the limit is a rational function of $h_{0}$, which is the desired claim.
@@
is well-defined, and rational in $h_{0}$ on $\mathbb{R}_{\geq 0}$, since all finite-dimensional averages taken with respect to the Laguerre ensembles therein can be reduced to finite linear combinations of ratios of $\Gamma$-functions, which simplify to rational functions. Combining this with the earlier claim, we see that under the given assumptions, both sides of (2) (modulo a common factor) define a rational function on $h_{0}\in[0,\infty)$. Moreover, by Theorem 1.3, these two functions agree on $h_{0}>\frac{1}{2}$, and hence, they must agree everywhere. This completes the proof of the desired result. ∎
@@
$\frac{1}{N^{2\sum_{m=2}^{k}mn_{m}}}\mathbb{E}_{N}^{(a,b)}\left[\mathrm{e}^{-\frac{t_{1}}{N^{2}}\sum_{j=1}^{N}\frac{1}{\mathsf{x}^{(N)}_{j}}}\prod_{m=2}^{k}\left(\sum_{j=1}^{N}(\mathsf{x}^{(N)}_{j})^{-m}\right)^{n_{m}}\right]$ $\xrightarrow[]{N\to\infty}\mathbb{E}\left[\mathrm{e}^{-t_{1}\mathfrak{e}_{1}(a)}\prod_{m=2}^{k}\left(\mathfrak{p}_{m}(a)\right)^{n_{m}}\right],$ (72)
Indeed, this follows from a combination of the distributional convergence in Proposition 2.5, Skorokhod’s representation theorem, and the uniform integrability of the sequence
$\left\{\mathrm{e}^{-\frac{t_{1}}{N^{2}}\mathfrak{p}_{1,N}^{(a,b)}}\prod_{m=2}^{k}(\mathfrak{p}_{m,N}^{(a,b)})^{n_{m}}\right\}_{N\geq 1},$
which is obtained by using Proposition 2.5 in combination with the elementary estimate $\mathrm{e}^{-x}\leq\frac{C_{k}}{x^{k}}$ that holds for $x\geq 0$, once $k$ is chosen so that the further parameter restriction in Proposition 2.5 holds. Similarly, we have that for any $m$ with $m<a+1$,
@@
Together with Theorem 1.9, if one defines the polynomials $\mathcal{P}_{m}^{(a)}(n_{2},\ldots,n_{k};t_{1})$ by,
$\mathcal{P}_{m}^{(a)}(n_{2},\ldots,n_{k};t_{1})=\lim_{N\to\infty}\frac{P_{m}^{(a,b)}(N;t_{1})}{N^{\sum_{q=2}^{k}2(q-1)n_{q}}},$ (74)
@@
for every $1\leq q\leq l$, (while doing the same for $\mathcal{G}_{N}^{(n_{2},\ldots,n_{k})}$), and establishes equivalent statements for $\mathcal{M}_{N,l,q}^{(n_{2},\ldots,n_{k})}$ as in Propositions 4.11 and 4.12 by matching the leading-order coefficients therein. Indeed, doing so and carefully carrying out the same inductive process as in the proof of Theorem 1.9, we find that the polynomials in $t_{1}$ appearing in the terms $\mathcal{M}_{N,l,q}^{(n_{2},\ldots,n_{k})}$ have the desired degrees as stated in Theorem 1.6.
The fact that the limit in (74) is independent of the parameter $b$ can be seen similarly based on the following observations. On one hand, we observe that in Proposition 4.11, the only term involving $b$ that appears in the expression of $\mathbf{H}_{l}^{(n_{2},\ldots,n_{k})}$ is the term $N^{2}\mathbf{P}_{2}^{(l)}\textbf{G}_{l-2}^{(n_{2},\ldots,n_{k})}$. Moreover, using the explicit expression for $\mathbf{P}_{2}^{(l)}$, it is easily seen that the highest-order contribution in $N$ coming from this particular term is the same as that of $N^{2}\mathbf{M}_{2}^{(l)}\textbf{G}_{l-2}^{(n_{2},\ldots,n_{k})}$, where $\mathbf{M}_{2}^{(l)}$ is given in the Appendix, and in particular, is independent of $b$. On the other hand, the only terms involving $b$ in Proposition 4.12 appear in the expressions for $\mathsf{F}_{0}^{(n_{2},\ldots,n_{k})}(t_{1})$ and $\mathsf{F}_{j}^{(n_{2},\ldots,n_{k})}(t_{1})$, $(j\geq 1)$, with the $b$-dependent terms contributing in a strictly smaller order of $N$. Hence, proceeding with the inductive process, where the inductive assumptions are in accordance with the formula (70), while keeping track of the coefficients of $N^{2l+\sum_{m=2}^{k}2(m-1)n_{m}}$ by means of these two observations, we get the desired conclusion regarding $b$-independence of the polynomials in (74). This concludes the proof of the theorem. ∎

### v2

@@ 4.2.3 A recursive relation for $\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{n,j}}$
where $P_{m,1}^{(a,-\frac{1}{2})}(t)$ and $P_{m,2}^{(a,-\frac{1}{2})}(t)$ are polynomials of $t$. Moreover, the coefficients of these polynomials are themselves polynomials in $N,a$. Note that, $\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\mathrm{e}^{-t\mathfrak{p}_{1,N}}\right]$ is $\sum_{k=1}^{m}kn_{k}$-times differentiable and it can be represented as $\exp\left(\int_{0}^{t}\frac{\sigma_{N}(t)}{t}\mathrm{d}t\right)$, where $\sigma_{N}(t)$ satisfies $\sigma$-Painlevé V equation (8). Hence, by examining the asymptotic expansion of $\sigma_{N}(t)$ in powers of $t$ at 0, up to $t^{\sum_{k=1}^{m}kn_{k}}$, we conclude that the coefficients of these terms can be obtained recursively and are rational functions of $a$, by an argument similar to that of [14, Section 4.2]). Moreover, this implies that the first $\sum_{k=1}^{m}kn_{k}+1$ Taylor coefficients of $\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\mathrm{e}^{-t\mathfrak{p}_{1,N}}\right]$ are of the form $\frac{\mathcal{Q}_{2}(N;a)}{\mathcal{Q}_{1}(a)}$, where $\mathcal{Q}_{1}(a)$ is a polynomial in $a$ and $\mathcal{Q}_{2}(N;a)$ is a polynomial in $a$, whose coefficients themselves are polynomials in $N$. After plugging this into the right-hand side of equation (4.4), while noting that the coefficients of the singular terms $\frac{1}{t^{l}}$ for $l=1,\dots,\sum_{k=1}^{m}kn_{k}$ vanish, since the value of the left-hand side of (4.4) is finite. We see that right-hand side of equation (4.4) must be a rational function in $a$ of the form described above; and therefore, so does the left-hand side of equation (4.4). Hence, by Newton’s formula, we readily conclude that $\mathbb{E}_{N}^{(a,-\frac{1}{2}{})}\left[\prod_{k=1}^{m}\mathfrak{e}_{N,k}^{h_{k}}\right]$ has the rational function form $\frac{Q_{2}(N;a)}{Q_{1}(a)}$, for certain polynomials $Q_{1}(a)$ and $Q_{2}(N;a)$ in $a$, whose coefficients of the powers of $a$ in $Q_{2}(N;a)$ are polynomials in $N$ whenever $\Re(a)>\sum_{k=1}^{m}kh_{k}-1$. On the other hand, the map $a\mapsto\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\prod_{k=1}^{m}\mathfrak{e}_{N,k}^{h_{k}}\right]$ is an analytic function by its definition on the set $\{a\in\mathbb{C}:\Re(a)>\sum_{k=1}^{m}h_{k}-1\}$. Thus, by the identity theorem, it must equal $\frac{Q_{2}(N;a)}{Q_{1}(a)}$ whenever $\Re(a)>\sum_{k=1}^{m}h_{k}-1$. Since $a=\sum_{k=0}^{m}h_{k}-\frac{1}{2}>\sum_{k=1}^{m}h_{k}-1$ when $h_{0}\geq 0$, $\frac{Q_{2}(N;a)}{Q_{1}(a)}$ is a rational function of $h_{0}$ on all of $h_{0}\geq 0$. Note that, when $h_{0}>\frac{1}{2}$, $N^{-2\sum_{k=1}^{m}{kh_{k}}}\frac{Q_{2}(N;a)}{Q_{1}(a)}$ converges as $N\rightarrow\infty$. Since this is a rational function, this convergence holds for any $h_{0}\geq 0$ and the limit is a rational function of $h_{0}$, which is the desired claim.
@@
is well-defined, and rational in $h_{0}$ on $\mathbb{R}_{\geq 0}$, since all finite-dimensional averages taken with respect to the Laguerre ensembles therein can be reduced to finite linear combinations of ratios of Gamma functions, which simplify to rational functions. Combining this with the earlier claim, we see that under the given assumptions, both sides of (2) (modulo a common factor) define a rational function on $h_{0}\in[0,\infty)$. Moreover, by Theorem 1.3, these two functions agree on $h_{0}>\frac{1}{2}$, and hence, they must agree everywhere. This completes the proof of the desired result. ∎
@@
$\frac{1}{N^{2\sum_{m=2}^{k}mn_{m}}}\mathbb{E}_{N}^{(a,b)}\left[\mathrm{e}^{-\frac{t_{1}}{N^{2}}\sum_{j=1}^{N}\frac{1}{\mathsf{x}^{(N)}_{j}}}\prod_{m=2}^{k}\left(\sum_{j=1}^{N}(\mathsf{x}^{(N)}_{j})^{-m}\right)^{n_{m}}\right]$ $\xrightarrow[]{N\to\infty}\mathbb{E}\left[\mathrm{e}^{-t_{1}\mathfrak{e}_{1}(a)}\prod_{m=2}^{k}\left(\mathfrak{p}_{m}(a)\right)^{n_{m}}\right].$ (75)
Indeed, an application of the elementary estimates $\mathrm{e}^{-x}\leq C_{r}x^{-r}$ for $x>0$, and $\sum_{j=1}^{N}y_{j}^{-l}\leq\left(\sum_{j=1}^{N}y_{j}^{-1}\right)^{l}$ for $l\in\mathbb{N}$ and $y_{1},\ldots,y_{N}>0$ gives that,
${\mathrm{e}^{-\frac{t_{1}}{N^{2}}\mathfrak{p}_{1,N}^{(a,b)}}\prod_{m=2}^{k}(\mathfrak{p}_{m,N}^{(a,b)})^{n_{m}}\leq\frac{N^{2r}C_{r}}{t_{1}^{r}(\mathfrak{p}_{1,N}^{(a,b)})^{r}}\prod_{m=2}^{k}(\mathfrak{p}_{m,N}^{(a,b)})^{n_{m}}\leq\frac{C_{r}N^{2r}}{t_{1}^{r}}(\mathfrak{p}_{1,N}^{(a,b)}})^{\sum_{m=2}^{k}mn_{m}-r}.$
Hence, picking an appropriate $r\in\mathbb{N}$ such that the parameter restriction in Proposition 2.5 holds, we conclude the the uniform integrability of the sequence
$\left\{N^{-2\sum_{m=2}^{k}mn_{m}}\mathrm{e}^{-\frac{t_{1}}{N^{2}}\mathfrak{p}_{1,N}^{(a,b)}}\prod_{m=2}^{k}(\mathfrak{p}_{m,N}^{(a,b)})^{n_{m}}\right\}_{N\geq 1}.$
Then, using the distributional convergence of the above sequence in Proposition 2.5 and Skorokhod’s representation theorem, we obtain (75). Similarly, we have that for any $m$ with $m<a+1$,
@@
Secondly, we claim that the convergence in (75) and (76) also hold at $t_{1}=0$. We only present here the argument for (75); which goes through mutatis mutandis for the convergence in (76). On one hand, by an argument similar to that of Corollary 1.10, for fixed b>1b>-1, the left-hand side of (75) at $t_{1}=0$ can be expressed as a rational function of $a$ and $N$ of the form $\frac{\tilde{Q}_{2}(N;a)}{\tilde{Q}_{1}(a)}$, for $\Re(a)>\sum_{m=2}^{k}mn_{m}-1$, where $\tilde{Q}_{1}(a)$ is a polynomial in $a$, and $\tilde{Q}_{2}(N;a)$ is a polynomial in $a$ whose coefficients are themselves polynomials in $N$. Meanwhile, the right-hand side of (75) at $t_{1}=0$ is also a rational function of $a$ for $a>\sum_{m=2}^{k}mn_{m}-1$. By Proposition 2.5, when $a>\sum_{m=2}^{k}mn_{m}$, the convergence of (75) at $t=0$ holds. Hence, by the properties of rational functions, this convergence extends to the full range $a>\sum_{m=2}^{k}mn_{m}-1$.
We now prove equation (6) of the theorem. Together with Theorem 1.9, if one defines the polynomials $\mathcal{P}_{m}^{(a)}(n_{2},\ldots,n_{k};t_{1})$ by,
$\mathcal{P}_{m}^{(a)}(n_{2},\ldots,n_{k};t_{1})=\lim_{N\to\infty}\frac{P_{m}^{(a,b)}(N;t_{1})}{N^{\sum_{m=2}^{k}2(m-1)n_{m}}},$ (77)
@@
for every $1\leq q\leq l$, (while doing the same for $\mathcal{G}_{N}^{(n_{2},\ldots,n_{k})}$), and establishes equivalent statements for $\mathcal{M}_{N,l,q}^{(n_{2},\ldots,n_{k})}$ as in Propositions 4.12 and 4.13 by matching the leading-order coefficients therein. Indeed, doing so and carefully carrying out the same inductive process as in the proof of Theorem 1.9, we find that the polynomials in $t_{1}$ appearing in the terms $\mathcal{M}_{N,l,q}^{(n_{2},\ldots,n_{k})}$ have the desired degrees as stated in Theorem 1.6.
The fact that the limit in (77) is independent of the parameter $b$ can be seen similarly based on the following observations. On one hand, we observe that in Proposition 4.12, the only term involving $b$ that appears in the expression of $\mathbf{H}_{l}^{(n_{2},\ldots,n_{k})}$ is the term $N^{2}\mathbf{P}_{2}^{(l)}\textbf{G}_{l-2}^{(n_{2},\ldots,n_{k})}$. Moreover, using the explicit expression for $\mathbf{P}_{2}^{(l)}$, it is easily seen that the highest-order contribution in $N$ coming from this particular term is the same as that of $N^{2}\mathbf{M}_{2}^{(l)}\textbf{G}_{l-2}^{(n_{2},\ldots,n_{k})}$, where $\mathbf{M}_{2}^{(l)}$ is given in the Appendix, and in particular, is independent of $b$. On the other hand, the only terms involving $b$ in Proposition 4.13 appear in the expressions for $\mathsf{F}_{0}^{(n_{2},\ldots,n_{k})}(t_{1})$ and $\mathsf{F}_{j}^{(n_{2},\ldots,n_{k})}(t_{1})$, $(j\geq 1)$, with the $b$-dependent terms contributing in a strictly smaller order of $N$. Hence, proceeding with the inductive process, where the inductive assumptions are in accordance with the formula (73), while keeping track of the coefficients of $N^{2l+\sum_{m=2}^{k}2(m-1)n_{m}}$ by means of these two observations, we arrive at the desired conclusion regarding $b$-independence of the polynomials in (77). This concludes the proof of the theorem. ∎